OpenStudy (anonymous):

(x + 3)2 = 18 extracting square root

OpenStudy (anonymous):

2x + 6 = 18 -6 -6 2x = 12 x = 6

OpenStudy (anonymous):

oh is it (x+3)^2 = 18 ? in that case it's x^2 + 3x +3x +9 = 18 x^2 +6x+9= 18 factoring (x+3)(x+3)=18 x+3 =18 -3 -3 x = 15 or x+3 = sq root of 18 which becomes x = sqrt18 -3

OpenStudy (anonymous):

sqrt of 18 is 3sqrt2

OpenStudy (anonymous):

this is extracting square root, the answer is a +-in a square root box

OpenStudy (anonymous):

Using the above rule, extract the square roots of the given quadratic equation. (x + 3)2 = 18 x + 3 = ± Correct: Your answer is correct. x + 3 = ± 3 Incorrect: Your answer is incorrect.

OpenStudy (anonymous):

well the answer can be plus or minus depening on which number plugged back into the equation gives you a true statement

OpenStudy (anonymous):

my first answer was square root of 18

OpenStudy (anonymous):

so the answer is supposed to be x= +/- sqrt18 (3sqrt2) -3 ? which should be positive ?

OpenStudy (anonymous):

im not sure what the boxes mean?

OpenStudy (anonymous):

it is the square root box it just wouldn't show sorry

OpenStudy (anonymous):

i believe it's just asking for the x+3?

OpenStudy (anonymous):

sorry i couldn't be more help. gotta go.

OpenStudy (anonymous):

thanks you have been of great help i appreciate that.

OpenStudy (triciaal):

extracting square roots